# Repoformer: Selective Retrieval for Repository-Level Code Completion

### v1

@@ 4.2 Evaluation Setup
We evaluate on RepoEval (Zhang et al., 2023), which consists of line, API, and function completion tasks created from 32 Python repositories. To improve repository coverage, we additionally leverage 1500 raw Python repositories from CrossCodeEval (Ding et al., 2023) to create a new chunk and function completion benchmark, which we call CCEval (details in Appendix D). To investigate the generalization to other languages, we also evaluated the original CrossCodeEval, which covers four languages: Python, Java, C#, and TypeScript (Section E.2).
@@
where $Lev$ is the Levenshtein distance (Levenshtein et al., 1966). We report $ES\times 100$ in all the tables following Zhang et al. (2023) for better readability. For execution-based evaluation, we report the unit test pass rate (UT). $\hat{Y}$ is said to pass the unit tests if replacing $Y$ with $\hat{Y}$ does not cause any unit test to fail.
@@
We experiment on two families of strong code LMs. CodeGen-Mono (Nijkamp et al., 2022a) is pretrained sequentially in natural language, multilingual code, and a Python corpus. StarCoder and StarCoderBase (Li et al., 2023b) are trained with fill-in-the-middle ability on a large corpus of multilingual code, GitHub issues, Git commits, and Jupyter notebooks. StarCoder is obtained by training StarCoderBase on an additional Python corpus.
@@ 5 Results
In this section, we first show that retrieval augmentation is often unhelpful for RAG with black-box code LMs. Then we demonstrate that our framework allows the model to avoid unnecessary retrievals and be more robust to the retrieved contexts, resulting in better accuracy and latency.
@@ 5.1 Is retrieval always helpful?
Table 1: The performance change on RepoEval function completion exhibited by four models from retrieved cross-file contexts. For the majority of the instances, RAG does not improve the performance. $\uparrow$”, ”$=$”, $\downarrow$” denote the counts for performance increase, no performance change, and performance drop.
RepoEval CCEval
Line API Function Chunk Function
Model Size Selective Policy EM ES EM ES UT ES EM ES ES
No Retrieval
1B - 43.44 67.77 37.81 66.54 22.20 47.65 31.08 60.09 47.49
3B - 49.00 72.12 40.44 69.02 24.84 51.22 36.14 64.65 49.88
StarCoderBase 7B - 51.88 74.03 43.31 70.79 25.49 52.28 38.88 66.61 52.45
StarCoder 16B - 55.25 76.07 44.50 71.00 34.73 53.60 42.58 69.40 54.20
Invariable Retrieval
1B - 51.19 72.30 43.94 69.17 25.71 55.64 37.22 63.73 50.50
3B - 56.69 76.68 47.00 72.62 29.67 57.68 42.26 67.74 53.39
StarCoderBase 7B - 59.44 78.15 49.56 73.65 31.43 58.51 44.44 69.53 55.41
StarCoder 16B - 61.25 79.24 51.12 74.50 42.86 60.96 47.90 71.90 58.06
Selective Retrieval
Repoformer 1B self-selection 51.90 74.50 43.50 71.00 24.00 53.10 38.52 68.08 52.09
$P$(<cc>) 54.40 76.00 46.10 72.70 28.79 57.30 41.92 69.97 53.71
3B self-selection 56.30 77.60 46.10 73.60 28.57 54.70 42.06 70.70 54.47
$P$(<cc>) 59.63 79.02 49.31 74.96 32.96 60.56 46.66 72.23 56.24
Table 2: Experiment results on RepoEval and CCEval. The best performance is underlined, and the best among models under 10B is boldfaced. Compared to StarCoderBase of the same size, Repoformer exhibits a strong performance, with Repoformer-3B outperforming other invariable retrieval models under 10B in most of the tasks. More importantly, Repoformer consumes a lower retrieval budget via selective retrieval. Among the two selective policies, $P$(<cc>) enables the best selective RAG performance.
In Figure 3 and Table 1, we evaluate four code LMs on API completion and function completion from RepoEval. For each model, we report the instance-level performance change from code completion only using $X_{l}$ and $X_{r}$ to retrieval-augmented code completion with $X_{l}$, $X_{r}$, and $CC$ (detailed prompts in Appendix A). The results reveal an intriguing pattern: retrieval improves LMs’ performance on only 20% or fewer instances. For the rest 80%, retrieval augmentation mostly does not affect the performance, but also actually harms the performance almost as often as it helps. The observed trends are consistent for both API and function completion and for both small-sized LMs ($\leq 2B$) and moderate-to-large LMs ($16B$). The generality of this observation is further confirmed by an analysis of Repoformer’s training data (Appendix D). Together, these findings highlight the suboptimality of the “invariable retrieval” approach and justify our selective retrieval proposal.
Next, we evaluate the code completion performance of Repoformer. We compare the following three settings [^3] We do not consider iterative retrieval because we found the first iteration contributes the majority of the performance gains.. For the first two baselines, we use the state-of-the-art single-iteration prompting pipeline (Zhang et al. (2023), detailed in Appendix A). We use StarCoder models due to their strong performance compared to the CodeGen family.
@@ 5.2 Repoformer achieves strong code completion performance via selective RAG
    Invariable Retrieval. This baseline always augments $X_{l}$ and $X_{r}$ with the retrieved $CC$.
@@
    • Self-selection. Retrieval is performed if <cc> is the most likely token following <eof>.
    • $P$(<cc>). If the probability of <cc> following <eof> is greater than a threshold $T$, retrieval augmentation is performed for this instance [^4] We find that $T=0.15$ for function completion and $T=0.2$ for the other tasks generally work well for $P$(<cc>). These two thresholds are always used unless otherwise stated..
The results are summarized in Table 2. Compared to no retrieval and invariable retrieval with StarCoderBase of the same size, Repoformer’s selective retrieval strategy exhibits strong performance improvements. Via the $P$(<cc>) strategy, Repoformer-3B can outperform StarCoderBase-7B on most of the tasks and metrics except EM for API completion, even outperforming the 5x larger StarCoder in terms of ES for API and chunk completion. We also show that the performance improvement from our paradigm can generalize to three languages beyond Python (Section E.2), the Repoformer-7B (Section E.2), as well as dense retrieval instead of Jaccard similarity (Section E.3). In later sections, we demonstrate that the observed success is due to an organic combination of the learned abstention ability and the improved robustness to retrieval.
In terms of code completion accuracy, the threshold-controlled $P$(<cc>) strategy outperforms the self-selection strategy on all the tasks. In the next section, we show that the two strategies represent different ways to trade-off between accuracy and inference budget.
@@ 5.3 Repoformer improves inference efficiency
We illustrate the benefits of Repoformer for saving the inference latency in a realistic “online serving” scenario.
@@
Selective Policy API Completion Line Completion
@@
invariable retrieval 72.02 $100\%$ $0\%$ 75.91 $100\%$ 0%
self-selection 71.04 $18\%$ $69\%$ 74.50 $19\%$ $61\%$
1B $P$(<cc>) 72.72 $61\%$ $28\%$ 76.00 $62\%$ $27\%$
invariable retrieval 74.66 $100\%$ $0\%$ 78.68 $100\%$ 0%
self-selection 73.60 $19\%$ $46\%$ 77.60 $20\%$ $43\%$
3B $P$(<cc>) 74.96 $78\%$ $17\%$ 79.02 $74\%$ $16\%$
Table 3: RAG latency of Repoformer with two self-selective RAG paradigms. %RAG = ratio of instances where RAG is performed. SU = Speedup compared to invariable retrieval (the higher, the better). Compared to the invariable retrieval baseline, the $P$(<cc>) strategy consistently demonstrates gains in both accuracy and latency. The self-selection strategy shows much larger latency gains with a small performance degradation.
Selective Policy API Completion Line Completion
@@
invariable retrieval 73.65 $0\%$ 78.15 $0\%$
7B Repoformer-1B 74.10 $24\%$ 78.31 $25\%$
invariable retrieval 74.50 $0\%$ 79.24 $0\%$
16B Repoformer-1B 74.84 $24\%$ 79.48 $24\%$
Table 4: Accuracy and latency of StarCoderBase-7B and StarCoder as the generation model and with Repoformer-1B as the policy model for selective RAG. SU = Speedup compared to invariable retrieval (the higher, the better). Compared to the invariable retrieval baseline, Repoformer’s selective decisions improve both the accuracy and latency of these larger LMs.
First, we consider $\mathcal{M}$ = Repoformer and present the results in Table 3. Line and API completion are presented to cover short and moderate target lengths [^5] We skip the function completion results as RepoEval uses very small repositories for function completion for easier unit testing.. Both selective strategies significantly improve the latency, with a different trade-off: using a fixed threshold for $P$(<cc>) results in improvements for both accuracy and latency compared to invariable retrieval, while using self-selection results in a larger latency gain with minor performance degradation (around 1.0 ES). It’s worth mentioning that the speed improvement from selective RAG could be further enhanced with a more advanced retrieval setup. For instance, dense retrieval on large repositories often consumes more than 80% of the entire RAG pipeline’s latency. In that case, a 20% RAG policy translates into more than 70% speedup. We empirically verify this statement in Section E.3.
Next, we consider using larger LMs as $\mathcal{M}$ in the framework and using $P$(<cc>) with Repoformer-1B as a plug-and-play selective RAG policy. As shown in Table 4, the selective predictions from Repoformer successfully reduce the latency of RAG with these 7x and 16x larger LMs by approximately 25% while improving their accuracy. The findings collectively indicate that Repoformer has acquired robust selective retrieval capabilities, which could partially transfer to other larger code LMs.
@@ 6 Analysis
Figure 4: An analysis of the instances where Repoformer-1B abstains from retrieval. We divide the instances into (1) the model answering correctly without retrieval (blue), the model making a mistake that cannot be improved by retrieval (yellow), and the model achieving better performance when retrieval is performed (red). The precision of abstention is over 0.8 on all tasks except for Function (RepoEval), which has a precision of 0.78.
@@
In Figure 4, we evaluate the precision of retrieval abstention decisions made by Repoformer’s $P$(<cc>) strategy. We find that the abstentions are accurate for over 80% instances across all the tasks. We also evaluate the calibration of the selective decisions and find Repoformer generally making near-calibrated predictions for line and API completion while the calibration is suboptimal for function completion with UT employed as the metric (Section E.1). We hypothesize that this could be caused by using ES to create the training signal and encourage future work to devise methods for labeling the quality of function completion more effectively.
@@
In Figure 5, we show the performance change caused by $CC$ on the instances where Repoformer requests for retrieval. Compared to StarCoderBase, Repoformer exhibits more and greater performance gains upon observing $CC$. The number of performance decreases is also significantly reduced, indicating an improved robustness.
Is Repoformer sensitive to threshold settings?
In Figure 1 (b), we present the code completion accuracy and latency of Repoformer as a function of the threshold. As the threshold increases, the model’s code completion performance first increases due to avoiding potentially harmful retrievals. At threshold 0.4, the model still maintains similar performance compared to invariable retrieval, with latency reduced by 50%. This result demonstrates that Repoformer can accommodate various threshold settings and provide a good accuracy-latency trade-off. We provide the visualization for other tasks in Section E.4.
@@
We fine-tune StarCoderBase-1B with the same setup as Repoformer and present the results on CCEval in Table 5. Although A1 has slightly better RAG performance, it fails to make meaningful selective decisions due to $\mathcal{L}_{eval}$ being outweighed by $\mathcal{L}_{gen}$ in long sequences: $P$(<cc>) is almost always 1. For A2, we find it only slightly outperforms Repoformer, suggesting learning $\mathcal{L}_{eval}$ does not harm the RAG ability a lot while bringing in the strong selective retrieval ability, which in turn boosts both accuracy and latency. A3 has the same performance for in-file completion as Repoformer, but exhibits worse RAG performance, indicating the necessity of training with $CC$. Finally, A4 achieves reasonable chunk completion performance but performs much worse in function completion. We hypothesize that placing $CC$ within the infilling part breaks the fill-in-the-middle semantics learned in StarCoder pre-training.
Model Selective Policy Chunk Completion Function Completion
@@
SC - - 0% 60.09 - 0% 47.49
- - 100% 63.73 - 100% 50.50
RF - - 0% 66.22 - 0% 49.77
$P$(<cc>) 0.20 75% 69.97 0.15 76% 53.71
- - 100% 69.95 - 100% 53.56
A1 - - 0% 66.14 - 0% 49.25
$P$(<cc>) 0.99 100% 70.21 0.99 100% 53.93
- - 100% 70.21 - 100% 53.93
A2 - - 0% 66.49 - 0% 49.02
- - 100% 70.45 - 100% 53.90
A3 - - 0% 66.25 - 0% 49.01
- - 100% 68.85 - 100% 52.12
A4 - - 0% 64.96 - 0% 25.44
$P$(<cc>) 0.10 86% 69.35 0.10 83% 26.50
- - 100% 69.19 - 100% 26.35
Table 5: Ablation study results. We report the performance on two tasks from the CCEval dataset. SC = StarCoderBase-1B. RF = Repoformer-1B. T = threshold for the retrieval policy. We found T = 0.10 works better for A4 and thus applied it to all the A4 results. %RAG = ratio of instances where RAG is performed.

### v2

@@ 4.2 Evaluation Setup
We evaluate on RepoEval (Zhang et al., 2023), which consists of line, API, and function completion tasks created from 32 Python repositories. To investigate the generalization to other languages, we also evaluated the original CrossCodeEval (Ding et al., 2023), which features line completion instances covering four languages: Python, Java, C#, and TypeScript (Section E.2). Observing that RepoEval has a limited repositrory coverage and that CrossCodeEval has a limited task coverage, we additionally leverage 1500 raw Python repositories from CrossCodeEval to create a new chunk and function completion benchmark, which we call CrossCodeLongEval. We detail the dataset creation process and basic statistics in Appendix D. For the rest of this paper, we will use CCEval to refer to both CrossCodeEval and CrossCodeLongEval interchangeably, and use the specific language and task (line, chunk, or function completion) to differentiate them.
@@
where $Lev$ is the Levenshtein distance (Levenshtein et al., 1966). We report $ES\times 100$ in all the tables following Zhang et al. (2023) for better readability. For execution-based evaluation, we report the unit test pass rate (UT). $\hat{Y}$ is said to pass the unit tests if replacing $Y$ with $\hat{Y}$ does not cause any unit test to fail. We implement simple post-processing procedures to handle common cases such as excessive lines in model’s outputs, which are documented in Appendix A.
@@
We experiment on two families of strong code LMs. CodeGen-Mono (Nijkamp et al., 2023b) is pretrained sequentially in natural language, multilingual code, and a Python corpus. StarCoder and StarCoderBase (Li et al., 2023b) are trained with fill-in-the-middle ability on a large corpus of multilingual code, GitHub issues, Git commits, and Jupyter notebooks. StarCoder is obtained by training StarCoderBase on an additional Python corpus.
@@ 5.1 Is retrieval always helpful?
In Table 1 and Figure 3, we evaluate four code LMs on function completion and API completion from RepoEval. For each model, we report the instance-level performance change from code completion only using $X_{l}$ and $X_{r}$ to retrieval-augmented code completion with $X_{l}$, $X_{r}$, and $CC$ (detailed prompts in Appendix A).
The results reveal an intriguing pattern: for repository-level code completion, the help from cross retrieval is often sparse. Specifically, retrieval improves LMs’ performance on only 20% or fewer instances. For more than 60% of the instances, retrieval augmentation does not affect the performance at all [^4] Upon a manual inspection, we find that most of the outputs in this category are also not changed by retrieval at all.. Finally, another 20% retrievals actually harm the performance, almost as often as the first case. The observed trends are consistent for both API and function completion and hold for both small-sized (1B and 2B) and moderate-to-large (around 16B) code LMs. The generality of this observation is further confirmed by an analysis of Repoformer’s training data, where we find that retrieval improves the performance for only fewer than 30% instances (Appendix D). Together, these findings highlight the suboptimality of the always retrieving and augmenting the cross-file contexts and thus motivate our selective retrieval proposal.
@@ 5.2 Repoformer achieves strong code completion performance via selective RAG
Table 1: The performance change on RepoEval function completion exhibited by four models from retrieved cross-file contexts. For the majority of the instances, RAG does not improve the performance. $\uparrow$”, “$=$”, $\downarrow$” denote the counts for performance increase, no performance change, and performance drop.
RepoEval CrossCodeLongEval
(Line) (API) (Function) (Chunk) (Function)
Size Model RAG Policy EM ES EM ES UT ES EM ES ES
1B StarCoderBase No 43.44 67.77 37.81 66.54 22.20 47.65 31.08 60.09 47.49
Always 51.19 72.30 43.94 69.17 25.71 55.64 37.22 63.73 50.50
Repoformer SelectiveG 51.90 74.50 43.50 71.00 24.00 53.10 38.52 68.08 52.09
SelectiveT 54.40 76.00 46.10 72.70 28.79 57.30 41.92 69.97 53.71
3B StarCoderBase No 49.00 72.12 40.44 69.02 24.84 51.22 36.14 64.65 49.88
Always 56.69 76.68 47.00 72.62 29.67 57.68 42.26 67.74 53.39
Repoformer SelectiveG 56.30 77.60 46.10 73.60 28.57 54.70 42.06 70.70 54.47
SelectiveT 59.63 79.02 49.31 74.96 32.96 60.56 46.66 72.23 56.24
7B StarCoderBase No 51.88 74.03 43.31 70.79 25.49 52.28 38.88 66.61 52.45
Always 59.44 78.15 49.56 73.65 31.43 58.51 44.44 69.53 55.41
Repoformer SelectiveG 56.00 76.63 48.06 75.03 30.77 55.27 43.80 72.46 56.14
SelectiveT 59.63 78.63 50.87 76.89 35.16 60.64 46.88 74.20 57.18
16B StarCoder No 55.25 76.07 44.50 71.00 34.73 53.60 42.58 69.40 54.20
Always 61.25 79.24 51.12 74.50 42.86 60.96 47.90 71.90 58.06
Repoformer SelectiveG 58.13 78.81 48.69 76.23 42.42 58.42 45.00 73.36 57.71
SelectiveT 61.75 80.34 51.88 77.93 44.18 62.58 49.18 75.50 58.93
Table 2: Experiment results on RepoEval and CrossCodeLongEval. The best performance among each model size is boldfaced. We use SelectiveG and SelectiveT to denote the greedy selection and the threshold selection strategy for selective retrieval. Repoformer greatly outperforms StarCoderBase of the same size while consuming a smaller retrieval budget. Among the two selective policies, threshold selection enables the best selective RAG performance.
Next, we evaluate the code completion performance of Repoformer. We compare the following three settings [^5] We do not consider iterative retrieval because we find that single-iteration RAG already achieves the majority of the performance gains from multi-iteration RAG.. For the first two baselines, we use the state-of-the-art single-iteration prompting pipeline (Zhang et al. (2023), detailed in Appendix A). We use StarCoder models due to their strong performance among the open-source code LMs.
@@
    Always Retrieving. This baseline always augments $X_{l}$ and $X_{r}$ with the retrieved $CC$.
@@
    • Greedy Selection. Retrieval is performed if <cc> is the most likely token following <eof>.
    • Threshold Selection. If the probability of <cc> following <eof> is greater than a threshold $T$, retrieval augmentation is performed [^6] We find that $T=0.15$ for function completion and $T=0.2$ for the other tasks generally work well. These two thresholds are always used unless otherwise stated..
The results are summarized in Table 2. Compared to no retrieval and always retrieving with StarCoderBase of the same size, Repoformer’s selective retrieval strategy exhibits strong performance improvements across all the tasks and both lexical-based and execution-based metrics. Via the threshold selection strategy, Repoformer-3B can outperform StarCoderBase-7B on most of the tasks and metrics except EM for API completion, even outperforming the 5x larger StarCoder in terms of ES for API and chunk completion. Finally, The Repoformer-16B model outperforms the strongest StarCoder baseline by 3%, averaged across all tasks, setting up the new start-of-the-art for repository-level code completion. We also experimentally confirm that the performance improvement from our framework can generalize to three languages beyond Python (Section E.2) as well as dense retrieval instead of Jaccard similarity (Section E.3). In later sections, we demonstrate that the observed success is due to both the ability to accurately abstain from retrieval and the improved robustness to retrieval.
In terms of code completion accuracy, the threshold selection strategy outperforms the greedy selection strategy on all the tasks. In the next section, we show that the two strategies represent different ways to achieve a good balance between accuracy and inference budget.
@@ 5.3 Repoformer improves inference efficiency
We illustrate the benefits of Repoformer for saving the inference latency in a realistic “online serving” setting.
@@
RAG Policy API Completion Line Completion
@@
Always 72.02 $100\%$ $0\%$ 75.91 $100\%$ 0%
SelectiveG 71.04 $18\%$ $69\%$ 74.50 $19\%$ $61\%$
1B SelectiveT 72.72 $61\%$ $28\%$ 76.00 $62\%$ $27\%$
Always 74.66 $100\%$ $0\%$ 78.68 $100\%$ 0%
SelectiveG 73.60 $19\%$ $46\%$ 77.60 $20\%$ $43\%$
3B SelectiveT 74.96 $78\%$ $17\%$ 79.02 $74\%$ $16\%$
Table 3: RAG latency of Repoformer with two self-selective RAG paradigms. %RAG = ratio of instances where RAG is performed. SU = Speedup compared to always retrieving (the higher, the better). Compared to the always retrieving baseline, the threshold selection strategy consistently demonstrates gains in both accuracy and latency. The greedy selection strategy shows much larger latency gains with a small performance degradation.
First, we consider $\mathcal{M}$ = Repoformer and present the results in Table 3. Line and API completion are presented to cover short and moderate target lengths [^7] We omit the function completion results as RepoEval uses very small repositories for function completion for easier unit testing.. Both selective strategies significantly improve the latency, with a different trade-off: threshold selection results in improvements for both accuracy and latency compared to always retrieving, while using greedy selection results in a larger latency gain with a minor performance degradation (around 1.0 ES). It is worth mentioning that the latency improvement from selective RAG could be further enhanced with a more advanced retrieval setup. For instance, conducting dense retrieval on large repositories often consumes more than 80% of the entire RAG pipeline’s latency. Then, a 20% RAG policy could translate into more than 70% speedup. We empirically verify this statement in Section E.3.
Next, we consider using diverse larger LMs as $\mathcal{M}$ in the code completion framework and using selectionT with Repoformer-1B as a plug-and-play selective RAG policy to decide whether retrieval should be performed. We experiment on a diverse set of LMs: StarCoderBase, Code Llama (Roziere et al., 2023) [^8] We accessed the model through Amazon SageMaker (https://docs.aws.amazon.com/sagemaker/)., CodeGen25 (Nijkamp et al., 2023a), and ChatGPT [^9] We use gpt-3.5-turbo-0613 via the OpenAI API.. As shown in Table 4, the selective predictions from Repoformer-1B successfully reduce the inference latency with different larger LMs by approximately 25% while improving their accuracy. Collectively, the findings indicate that Repoformer has acquired robust selective retrieval capabilities that could generalize to diverse types of code LMs.
Model RAG Policy API Completion Line Completion
@@
Always Retrieving 73.65 $0\%$ 78.15 $0\%$
SCB-7B Repoformer-1B 74.10 $24\%$ 78.31 $25\%$
Always Retrieving 74.50 $0\%$ 79.24 $0\%$
SCB-16B Repoformer-1B 74.84 $24\%$ 79.48 $24\%$
Always Retrieving 63.07 0% 68.42 0%
CG25-7B Repoformer-1B 63.37 20% 68.86 29%
Always Retrieving 58.75 0% 59.99 0%
CL-7B Repoformer-1B 58.91 25% 60.47 28%
Always Retrieving 61.08 0% 61.58 0%
CL-16B Repoformer-1B 62.10 32% 62.45 30%
Always Retrieving 63.38 0% 61.76 0%
ChatGPT Repoformer-1B 64.01 28% 61.92 18%
Table 4: Accuracy and latency of larger code LMs as the generation model and with Repoformer-1B as the policy model for selective RAG. SCB = StarCoderBase, CG25 = CodeGen25, CL = Code Llama. SU = Speedup compared to Always Retrieving (the higher, the better). Compared to the Always Retrieving baseline, Repoformer’s selective decisions improve both the accuracy and latency of these larger LMs.
@@ 6 Analysis
Is Repoformer sensitive to threshold settings?
In Figure 4, we present the code completion accuracy and latency of Repoformer as a function of the threshold. As the threshold increases, the model’s code completion performance first increases due to avoiding potentially harmful retrievals. At threshold 0.4, the model still maintains similar performance compared to always retrieving, with latency reduced by 50%. This result demonstrates that Repoformer can accommodate various threshold settings and provide a good accuracy-latency trade-off. We provide the visualization for other tasks in Section E.4.
@@
In Figure 5, we evaluate the precision of retrieval abstention decisions made by Repoformer’s threshold selection strategy. We find that the abstentions are accurate for over 80% instances across all the tasks: when Repoformer abstains from retrieval, its code completion prediction either is already correct without retrieval or cannot be improved by retrieval. We also evaluate the calibration of the selective decisions and find Repoformer generally making near-calibrated predictions for line and API completion while the calibration is suboptimal for function completion with UT employed as the metric (Section E.1). We hypothesize that this could be caused by using ES to create the training signal and encourage future work to devise methods for labeling the quality of function completion more effectively.
Figure 4: The accuracy and latency change with different threshold settings. Selective Retrieval with Repoformer achieves better accuracy and better latency than always retrieving. In addition, this behavior is relatively insensitive to the threshold.
Figure 5: An analysis of the instances where Repoformer-1B abstains from retrieval. We divide the instances into (1) the model answering correctly without retrieval (dark blue), the model making a mistake that cannot be improved by retrieval (light blue), and the model achieving better performance when retrieval is performed (red). The precision of abstention is over 0.8 on all tasks except for Function (RepoEval), which has a precision of 0.78.
@@
In Figure 6, we show the performance change caused by $CC$ on the instances where Repoformer requests for retrieval. Compared to StarCoderBase, Repoformer exhibits more and greater performance gains upon observing $CC$. The number of performance decreases is also significantly reduced, indicating an improved robustness to the potentially irrelevant retrieval contexts. In Table 8 in the appendix, we further study the effect of using dense retrieval. Although dense retrieval returns an arguably different context distribution compared to sparse retrieval, Repoformer still exhibits strong improvements in both quality and latency.
@@
We fine-tune StarCoderBase-1B with the same setup as Repoformer and present the results on CCEval in Table 5. Although A1 has slightly better RAG performance, it fails to make meaningful selective decisions due to $\mathcal{L}_{eval}$ being outweighed by $\mathcal{L}_{gen}$ in long sequences: the probability of <cc> is almost always 1. For A2, we find it only slightly outperforms Repoformer, suggesting learning $\mathcal{L}_{eval}$ does not harm the RAG ability a lot while bringing in the strong selective retrieval ability, which in turn boosts both accuracy and latency. A3 has the same performance for in-file completion as Repoformer, but exhibits worse RAG performance, indicating the necessity of training with $CC$. Finally, A4 achieves reasonable chunk completion performance but performs much worse in function completion. We hypothesize that placing $CC$ within the infilling part is detrimental due to breaking the fill-in-the-middle semantics learned in StarCoder pre-training.
Model RAG Policy Chunk Completion Function Completion
@@
SC No - 0% 60.09 - 0% 47.49
Always - 100% 63.73 - 100% 50.50
RF No - 0% 66.22 - 0% 49.77
SelectiveT 0.20 75% 69.97 0.15 76% 53.71
Always - 100% 69.95 - 100% 53.56
A1 No - 0% 66.14 - 0% 49.25
SelectiveT 0.99 100% 70.21 0.99 100% 53.93
Always - 100% 70.21 - 100% 53.93
A2 No - 0% 66.49 - 0% 49.02
Always - 100% 70.45 - 100% 53.90
A3 No - 0% 66.25 - 0% 49.01
Always - 100% 68.85 - 100% 52.12
A4 No - 0% 64.96 - 0% 25.44
SelectiveT 0.10 86% 69.35 0.10 83% 26.50
Always - 100% 69.19 - 100% 26.35
Table 5: Ablation study results. We report the performance on two tasks from the CCEval dataset. SC = StarCoderBase-1B. RF = Repoformer-1B. T = threshold for the SelectiveT policy. We found T = 0.10 works better for A4 and thus applied it to all the A4 results. %RAG = ratio of instances where RAG is performed.